OpenStudy (anonymous):

Let f(x) = ln(x^7) and g(x) = (ln(x))^7. Then (df)/(dx) =

OpenStudy (campbell_st):

for logs the derivative dy/dx = f'(x)/f(x) f(x) = x^7 f'(x) = 7x^6 df(x)/dx = 7x^6/x^7 = 7/x

OpenStudy (anonymous):

what if its (df)/(dx)|_(x = 1)?